Give an example of a divergent sequence which has two convergent sequences. Justify

your claim.


1
Expert's answer
2021-04-29T17:35:34-0400

Example of a divergent sequence which has two convergent sequences.


Let's consider the sequences "x_n = \begin{cases}
 0 &\text{if } n-odd \\
 1 &\text{if } n-even
\end{cases}" and "y_n = \begin{cases}
 1 &\text{if } n-odd \\
 0 &\text{if } n-even
\end{cases}"



Since the subsequences "x_{2k}=1_{k\to\infty}\to1" and "x_{2k-1}=0_{k\to\infty}\to0" have different limit points, then the sequence "x_n = \begin{cases}
 0 &\text{if } n-odd \\
 1 &\text{if } n-even
\end{cases}" is divergent.


Since the subsequences "y_{2k}=1_{k\to\infty}\to1" and "y_{2k-1}=0_{k\to\infty}\to0" have different limit points, then the sequence "y_n = \begin{cases}
 1 &\text{if } n-odd \\
 0 &\text{if } n-even
\end{cases}" is divergent.


Now, we can consider the sum of those sequences:

"x_n+y_n = \begin{cases}
 1 &\text{if } n-odd \\
 1 &\text{if } n-even
\end{cases}= 1_{n\to \infty}\to1," Hence the sequence "(x_n+y_n)" is convergent.


And also, we can consider the product of those sequences:

"x_n\times y_n = \begin{cases}

 0 &\text{if } n-odd \\

 0 &\text{if } n-even

\end{cases}= 1_{n\to \infty}\to1," Hence the sequence "(x_n\times y_n)" is convergent.


Final Answer: "x_n = \begin{cases}
 0 &\text{if } n-odd \\
 1 &\text{if } n-even
\end{cases}" , "y_n = \begin{cases}
 1 &\text{if } n-odd \\
 0 &\text{if } n-even
\end{cases}"